The length of a rectangle exceeds its width by 7 units and its area is 228 square units. What is the length of the rectangle?

Found 2 solutions by scott8148, 2332:
Answer by scott8148(6628)

L (L - 7) = 228

L^2 - 7L - 228 = 0

factoring ___ (L - 19)(L + 12) = 0

Answer by 2332(1)